OpenStudy (anonymous):

need help and show ur solution so i can understand it use implicit differentiation

OpenStudy (anonymous):

\[x(x+y)=x ^{4}y ^{2}-y\]

OpenStudy (amistre64):

ok, try to do what comes naturally .. what do you come up with?

OpenStudy (amistre64):

you know product rule and power rule .... those dont change

OpenStudy (anonymous):

plss show how u solve it . im confused in this problem

OpenStudy (amistre64):

i need to see your confusion to know where it resides. show me what you think should happen .... lets start with one piece of it, you see its a product .... what should we apply? x(x+y)

OpenStudy (anonymous):

\[x ^{2}+xy\]???

OpenStudy (anonymous):

i really dont know

OpenStudy (amistre64):

wasnt what i was going for, but that is a correct move nonetheless so, what is the d/dx of x^2?

OpenStudy (amistre64):

what do you normally do when you see x^2 ?

OpenStudy (anonymous):

2x

OpenStudy (amistre64):

good now, is that other term, xy, a product?

OpenStudy (anonymous):

no????

OpenStudy (amistre64):

xy means x times y and a product is math vocabulary for: "we multiply things together"

OpenStudy (amistre64):

xy is a product of x times y, therefore it IS a product what is the product rule for derivatives?

OpenStudy (anonymous):

xy

OpenStudy (amistre64):

spose we have to functions f(x) and g(x); or simply f and g the derivative of fg would simply be f'g + fg' does this ring a bell?

OpenStudy (anonymous):

xy + xy' ???????

OpenStudy (amistre64):

very close :) you prolly missed the x' mark by accident, but to conform this to what i said spose we have to functions x(x) and y(x); or simply x and y the derivative of xy would simply be x'y + xy' x' = dx/dx which is just 1 so we tend to exclude it x'y + xy' = y + xy'

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

how did u get y+xy'??

OpenStudy (amistre64):

i wrote it all out ... im not sure how much clearer i could have been with it. so far we got \[x(x+y)=x ^{4}y ^{2}-y\] \[x^2+xy=x ^{4}y ^{2}-y\] \[\frac d{dx}(x^2+xy=x ^{4}y ^{2}-y)\] \[\frac d{dx}(x^2)+\frac d{dx}(xy)=\frac d{dx}(x ^{4}y ^{2})-\frac d{dx}(y)\] \[\frac {dx}{dx}2x+\frac {dx}{dx}y+x\frac {dy}{dx}=\frac d{dx}(x ^{4}y ^{2})-\frac d{dx}(y)\] \[2x+y+xy'=\frac d{dx}(x ^{4}y ^{2})-\frac d{dx}(y)\] halfway there

OpenStudy (amistre64):

the d/dx (y) at the end of it should seem easy enough all thats left is the product of the powers

OpenStudy (amistre64):

you might notice that we havent learned any new derivative rules ....

OpenStudy (anonymous):

im trying to understand it

OpenStudy (amistre64):

lets try some side stuff; if i ask you what the derivative of x^3 is, what would you say it is?

OpenStudy (anonymous):

3x^2

OpenStudy (amistre64):

but your almost right, i never said to derive it "with respect to x" the derivative of x^3 is 3x^2 x' becasue of the chain rule

OpenStudy (amistre64):

what is the derivative of 2y^2?

OpenStudy (anonymous):

4y

OpenStudy (amistre64):

dont forget the chain rule, without knwoing what the derivative is "with respect to" you have to pop out the prime of the variable derivatove of 2y^2 = 4y y'

OpenStudy (amistre64):

what is the derivative of 6t ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

if the derivative is y?? you have to put the y'?

OpenStudy (amistre64):

your confusing a derivative with " a derivative with respect to __" without knowing the "with respect to" variable we simply dont know that value of what the chain rule pops out

OpenStudy (amistre64):

the derivative of y is: y' the derivative of r is: r' the derivative of p is: p' the derivative of B is: B' the derivative of 4g^3 is: 12g^2 g' the derivative of x^2y is: 2xx' y + x^2 y'

OpenStudy (amistre64):

you only begin to learn derivatives "with respect to x" but thats just to baby you into derivatives

OpenStudy (amistre64):

any of this making sense yet?

OpenStudy (amistre64):

if not, then youve simply learned derivatives the wrong way

OpenStudy (anonymous):

not yet